OpenStudy (jordanloveangel):

How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

Attach it

OpenStudy (anonymous):

What's the equation?

OpenStudy (jordanloveangel):

i think its C

OpenStudy (anonymous):

B Infinitely many, I think.

OpenStudy (anonymous):

Because the equation is true for all real numbers

OpenStudy (jordanloveangel):

but when i did it i got 0

OpenStudy (jordanloveangel):

Simplifying 3(4x + -3) + -7x = 5x + -9 Reorder the terms: 3(-3 + 4x) + -7x = 5x + -9 (-3 * 3 + 4x * 3) + -7x = 5x + -9 (-9 + 12x) + -7x = 5x + -9 Combine like terms: 12x + -7x = 5x -9 + 5x = 5x + -9 Reorder the terms: -9 + 5x = -9 + 5x Add '9' to each side of the equation. -9 + 9 + 5x = -9 + 9 + 5x Combine like terms: -9 + 9 = 0 0 + 5x = -9 + 9 + 5x 5x = -9 + 9 + 5x Combine like terms: -9 + 9 = 0 5x = 0 + 5x 5x = 5x Add '-5x' to each side of the equation. 5x + -5x = 5x + -5x Combine like terms: 5x + -5x = 0 0 = 5x + -5x Combine like terms: 5x + -5x = 0 0 = 0 Solving 0 = 0

OpenStudy (anonymous):

If 0=0 than it means the equation is true for any number plugged in.

OpenStudy (anonymous):

you didn't get "x=0" as a solution, you got "0=0"

OpenStudy (anonymous):

Do you kind of get it now?

OpenStudy (anonymous):

You did great work as well by the way

OpenStudy (jordanloveangel):

oh thanx so its b right

OpenStudy (anonymous):

Yeah man

OpenStudy (jordanloveangel):

thanx

OpenStudy (jordanloveangel):

can u help me with 2 more

OpenStudy (anonymous):

Uhm yeah i can try

OpenStudy (jordanloveangel):

Which expressions would complete this equation so that it has infinitely many solutions? 8 + 2(8x – 6) = Choose exactly two answers that are correct. A. 2(4x + 7) B. 9x – 10 C. 16x – 4 D. 4(4x – 1)

OpenStudy (jordanloveangel):

thanx

OpenStudy (anonymous):

Alright give me a minute.

OpenStudy (jordanloveangel):

kk

OpenStudy (anonymous):

The two are C. and D.